The healthcare provider has ordered Labetalol 1 mg/min for a patient with hypertension. The medication is delivered to the unit labeled 200mg/200mL 5% Dextrose. The patient weighs 86 kg.

The nurse should set the pump at what rate in mL per hour?
Answer #1

ANSWER.

* Calculation of rate(ml/hr).

         Rate (ml/hr) = Volume(ml) / Time(hr)

      Calculation of volume

            Volume(ml) = Desired dose(D)/Dose in hand (H) * Quantity(Q)

    Here

         Desired dose(D) = 1 mg/min.

         Dose in hand (H) = 200 mg.

               Quantity (Q) = 200 ml.

   As per above formula

           Volume(ml) of drug required = D/H * Q

                                                         = 1/200 * 200

                                                         = 0.005 * 200

                                                         = 1 ml/min

        Convert ml/min to ml/hr by multiply with 60

              So

                      ml/hr = 1 * 60

                               = 60 ml/hr.